A file system control block has the following address method: 10 direct block address, 10-bit single indirect address, and 10-bit double indirect address. The block size is 8K bytes. What address method will be used for the data files of 100K, 10M, and 10G bytes respectively? List the detailed computation to justify your answer.
Answer #1

Disk blocks= 8K bytes
10 direct block address
8K*10= 80K can be accessed directly

Single indirect block: –
10‐bit address, we have 1 byte per block pointer,
so 8KB/1 = 8192 blocks –
Can access file size: 8192 * 8 KB = 64 MB

Double‐indirect block: –
Can access file size: 8192 * 8192 * 8 KB = 512 GB

100K,10M single-indirect address method.

10G double-indirect address method